OpenStudy (anonymous):

The solution set for 4x^2-11x=10? Fractional form.

OpenStudy (anonymous):

I got \[11+\sqrt{287}/8 , 11-\sqrt{287}/8\]

OpenStudy (amistre64):

4x^2-11x=10 ; divide out the 4 x^2 - (11/4)x = 10/4 ; "complete the square", which is just half of (11/4), and ^2 it x^2 - (11/4)x + (11/8)^2 = 10/4 + (11/8)^2; clean it up (x - 11/8)^2 = 10/4 + 121/64 (x - 11/8)^2 = 10(16)/4(16) + 121/64 (x - 11/8)^2 = (160+121)/64 (x - 11/8)^2 = 281/64 ; now sqrt the sides x - (11/8) = +- sqrt(281)/8 ; and +11/8 x = (11/8) +- sqrt(281)/8 ; simplfy as wanted

OpenStudy (amistre64):

and of course dbl chk me math ;)